OpenStudy (anonymous):

Solve for y: x-2y over 5=1

OpenStudy (anonymous):

x-2y = 5, then -2y = 5-x y = (-5-x)/-2y y = (x+5)/2y

OpenStudy (anonymous):

oops, typo, should be y = (x+5)/2

OpenStudy (anonymous):

so it should be y = (5-x)/-2

OpenStudy (anonymous):

my last post was the answer.

OpenStudy (anonymous):

okay, so is this the 3 answer...?

OpenStudy (anonymous):

thank you!!! YOU are much help!